OpenStudy (anonymous):

What is the sum of the first 10 terms of the sequence defined by an=4n-3 A:37 B:190 C:380 D:384

OpenStudy (mathmale):

Dear Tamisha, The info contained in http://www.mathsisfun.com/algebra/sequences-sums-arithmetic.html may help you solve the problem you've posted. If not, continue to ask questions here. Good luck!

OpenStudy (anonymous):

a(n) = 4n-3 a(1)=4(1)-3= 4-3=1 a(2)=4(2)-3= 8-3 =5 a(3)=4(3)-3=12-3=9 So continue one plugging in terms up till 10 then add up all the #'s :D

OpenStudy (mathmale):

that's definitely one way in which this problem can be done, but unless you're using a spreadsheet like Excel, or plenty of elbow grease, it takes a while! There are summation formulas in http://www.mathsisfun.com/algebra/sequences-sums-arithmetic.html that may speed the process (if you're interested, that is).